9. A current source of 4 cos ar is connected across a 10 Ω resistor and 20 mH induc- tor in series. Find the phasor voltage and the real-time voltage across the series combination if the angular frequency is 377 rad/s. Ans. 50.1 exp(丿37.67, u(1) = 50.1 cos(377t + 37.67.

need to get 50.1exp(j37.6) and 50.1cos(377t+37.6)
Answer #1

Here inductive reactance Xl = 0.02×377= 7.54 ohm

So inpidence z = sqroot ( 10^2+7.54^2) =12.524 ohm

So magnitude of voltage V = i×z = 4×12.524 =50.1 V

Now phase angle Theta = atan(7.54/10) = 37°

So realtime voltage V = 50.1cos(377t +37)

And phasor voltage V = 50.1exp(j37)